Grade 9Physical Chemistry

a certain mass of a gas occupied 850ml at a pressure of 760mm of hg on increasing the pressure it was found that the volume of the gas was 75% of its initial value .assuming constant temperature find the final pressure of the gas ?

saipurvi
10 Years agoGrade 9

2 Answers

Ankit Jaiswal
10 Years ago
we know that for an ideal gas
pv=nrt
and hence 
v=nrt/p
in the question it is given that the same gas is kept afterwards at a lower pressure and hence the no. of moles is same 
so we know that the value of n,r,t are same (r is gas constant)
so 
\propto 1/p
we also knoe that the final volume of gas is 75% of initial volume
=850*75/100 ml
so 
volume initially(L)/volume final(L) =pressure final(atm) /pressure innitially(atm)
so
(850/1000) / (850*75/(100*1000))= (760/760) /p_{f}
so 
p_{f} =0.75 atm
which equals 570mm of hg
i hope this helps if so plz approve this ans
 
 
 
 
Vikas TU
ApprovedApproved Tutor Answer10 Years ago
T = constant
from Boyle’ law,
P1V1 = P2V2
Therfore,
760*850 = 0.75*850*P
P = 760 / 0.75
  = 1013.33 mm og Hg